OpenStudy (anonymous):

a The screams from the roller coaster were being measured that day. Our normal speaking sound level is measured 50 decibels (dB). The screams from ‘Leviathan’ were measured at 110 dB. How much more intense is the screams from the Leviathan than our normal sound level.

OpenStudy (kaiz122):

i think you just subtract 50 from 110. 110-50=?

OpenStudy (ganpat):

@kaiz122 : i guess, probably percentage may required..

OpenStudy (anonymous):

no. i have to use sound equation. but i dont know how

OpenStudy (ganpat):

oops, m not aware of that.. if u can provide eqtn from book.. than we might be able to help u...

OpenStudy (anonymous):

L= 10log(I/I\[L= 10 \log ( I/Io )\]

OpenStudy (ganpat):

so, L = 10 log (110/50) i guess, Io is normal decibel, while I is increased decibel.. can u calculate further.. You may use calculator..

OpenStudy (ganpat):

if m right, than it would be L = 7.88.. with its units...

OpenStudy (anonymous):

but some one said it is 1000000000 more instanse how do u get that

OpenStudy (ganpat):

wow... i have no idea.. sorry !!!

OpenStudy (anonymous):

no one help. i need it my today :(

OpenStudy (ganpat):

i wished i could.. ohkey, lets ask to some genius minds over here.. @ParthKohli , @waterineyes , @ganeshie8 ... Guys can anyone help, plz ??

OpenStudy (anonymous):

Sorry no clue.. @Ganpat

OpenStudy (anonymous):

thnakx for the help Ganpat its ok

ganeshie8 (ganeshie8):

50 = 10 log(I1/Io) ------- (1) 110 = 10 log(I2/Io) -------(2)

ganeshie8 (ganeshie8):

you need to take I2/I1 to know how intense is the I2 compared to I1

ganeshie8 (ganeshie8):

from (1), 50/10 = log(I1/Io) => 5 = log(I1/Io) => 10^5 = I1/Io => I1 = Io x 10^5 ------------------(3)

ganeshie8 (ganeshie8):

from (2), 110/10 = log(I2/Io) => 11 = log(I2/Io) => 10^11 = I2/Io => I2 = Io x 10^11 -----------------------(4)

ganeshie8 (ganeshie8):

combine (3) & (4) and take the ration I2/I1

OpenStudy (anonymous):

so lastly Ihow do u do

OpenStudy (anonymous):

you mean multiple 3 qnd 4

ganeshie8 (ganeshie8):

do (4) / (3) => \(\huge \frac{I2}{I1} = \frac{ Io . 10^{11}}{Io . 10^5} \) \(\huge = \frac{10^{11}}{10^5}\) \(\huge = 10^6\)

ganeshie8 (ganeshie8):

I2 is 10^6 times more powerful than I1

OpenStudy (anonymous):

thank youu soo much

ganeshie8 (ganeshie8):

yw =))